# xGen-MM (BLIP-3): A Family of Open Large Multimodal Models

### v2

@@ Abstract
This report introduces xGen-MM (also known as BLIP-3), a framework for developing Large Multimodal Models (LMMs). The framework comprises meticulously curated datasets, a training recipe, model architectures, and a resulting suite of LMMs. xGen-MM, short for xGen-MultiModal, expands the Salesforce xGen initiative on foundation AI models. Our models undergo rigorous evaluation across a range of tasks, including both single and multi-image benchmarks. Our pre-trained base model exhibits strong in-context learning capabilities and the instruction-tuned model demonstrates competitive performance among open-source LMMs with similar model sizes. In addition, we introduce a safety-tuned model with DPO, aiming to mitigate harmful behaviors such as hallucinations and improve safety. We open-source our models, curated large-scale datasets, and our fine-tuning codebase to facilitate further advancements in LMM research. Associated resources will be available on our project page above. Refer to caption Refer to caption (a) BLIP-2 framework (b) xGen-MM (BLIP-3) framework Figure 1: We introduce xGen-MM (BLIP-3), a framework (b) for developing Large Multimodal Models (LMMs). Our framework improves upon BLIP-2 (a) [1] by (1) increasing the richness, scale, and diversity of training data, (2) replacing the Q-Former layers with a more scalable vision token sampler, and (3) simplifying the training process via the unification of the training objectives to a single loss at every training stage. The resulting suite of LMMs can perform various visual language tasks and achieve competitive performance across benchmarks.
@@ 1 Introduction
Large Multimodal Models (LMMs) have attracted significant attention with their potential applications and emergent capabilities. Recent advancements in both proprietary models [2, 3, 4, 5] and open-source LMMs [6, 1, 7, 8, 9, 10, 11] highlight the rapid progress and growing interest in this field. However, despite these advancements, there is still a gap between open-source models and proprietary ones in terms of access to open weights, training recipes, and curated datasets. Such limitations hinder the open-source communities from replicating, understanding, and improving LMMs.
@@
Recent works have demonstrated that large-scale and high-quality data are essential for training robust LMMs [8, 9, 10, 11, 12]. BLIP-2 [1] was one of the pioneering efforts in exploring LMMs, which leveraged synthetic data to achieve impressive results at the time (Figure 1 (a)). However, the data used in BLIP-2 lacks the scale, quality, and diversity required to reach competitive performance compared to more modern LMMs nowadays. In addition, BLIP-2 employs an intricate Q-Former [1] architecture to bridge the vision and language modalities, coupled with a suite of complex training objectives (ITM, ITC, and ITG losses), both of which pose obstacles for larger-scale training. Moreover, BLIP-2 supports only single-image input, whereas interleaved multimodal data formats are the most natural form of multimodal data [13].
@@
In response to these challenges, we introduce xGen-MM (BLIP-3) (Figure 1 (b)), a new framework designed to scale up LMM training by utilizing an ensemble of multimodal interleaved datasets, curated caption datasets, and other publicly available datasets [14, 15, 16, 17]. xGen-MM, short for xGen-MultiModal, further expands our previous generative AI initiatives and corresponding foundation models for text xGen [18], code generation codeGen [19, 20], function calling APIGen [21], among others. In xGen-MM (BLIP-3), as illustrated in Figure 2, we streamline the model architecture by replacing the Q-Former [1] with a more scalable vision token sampler (specifically, a perceiver resampler [22]) and simplifying the training objectives to focus solely on the auto-regressive loss of text tokens in a multimodal context. Our primary focus is on dataset curation and scaling up the training data. Recently, our BLIP-3 team introduced two large-scale, high-quality datasets: MINT-1T [12], a trillion-token scale interleaved dataset; and BLIP3-KALE, a knowledge-augmented high-quality dense captions dataset. In this technical report, we introduce two additional specialized datasets: BLIP3-OCR-200M, a large-scale dataset with dense OCR annotations; and BLIP3-GROUNDING-50M, a large-scale visual grounding dataset.
@@
In addition to these datasets, we are committed to open-sourcing the series of models developed in this work, including the base, instruction-tuned, and DPO models. Along with the model release, we also provide code for easy fine-tuning of our base model on custom datasets. By making these resources publicly available, we aim to make LMM research and development more accessible to the community, and we encourage researchers and practitioners to use our models and datasets to understand and further explore the potential and emergent abilities of LMMs.
@@
Figure 2: Overview of the xGen-MM (BLIP-3) framework. Free-form interleaved images and texts from the ensembled interleaved and caption datasets are input into the framework, with each modality undergoing a separate tokenization process to be fed into the pre-trained LLM in natural order. A standard auto-regressive loss is then applied to the text tokens. The Vision Transformer is kept frozen during training, while all other parameters, including the token sampler and the pre-trained LLM, are trained.
@@ 2 Related Work
Recent advancements in Large Multimodal Models (LMMs) have explored two main architectural approaches: the cross-attention style [22, 23] and the decoder-only style. The cross-attention approach, exemplified by models like Flamingo [22, 23] and Llama 3.1 [5], integrates vision and language modalities through a complex attention mechanism to enable deep multimodal understanding. Another approach is the decoder-only architecture [1, 7, 8, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], which we adopt in xGen-MM (BLIP-3), offers a more streamlined solution. This approach connects pre-trained language models to visual inputs using lightweight connectors, simplifying the integration process while maintaining robust multimodal capabilities. The effectiveness of this architecture is evident in models such as MM1 [9], VILA [10], LLaVA [8], phi3-vision [37], and Otter [38].
Training methodologies for LMMs typically follow one of the two strategies. The first one uses a light pre-training procedure and heavily relies on visual instruction tuning, as seen in the LLaVA series [8, 29]. Extensive research has been conducted on creating effective instruction-tuning data for a variety of tasks [32, 39, 40, 41, 42, 43]. The second strategy involves extensive pre-training on large-scale, diverse datasets, followed by visual instruction fine-tuning. This approach, employed by models like MM1 and Idefics2 [11], infuses broad knowledge into the model, which is then fine-tuned to align with human-like interactions and safety standards. While MM1 [9] provides extensive ablations and studies on the recipes aimed at improving LMMs, it releases limited resources for practitioners to adopt the model (e.g., MM1 models and datasets are close-sourced). Idefics2 [11] is a more recent open-source work that open-sources a suite of models along with detailed training strategies and data recipes, but Idefics2 mostly resues existing datasets in their experiments without contributing new ones.
In this work, we present xGen-MM (BLIP-3). Unlike previous works, xGen-MM (BLIP-3) is an open-source family of a series of models, data recipes, fine-tuning code, and two large-scale foundational multimodal datasets, which we hope will enable and advance future research in this area.
@@ 3 Model Architecture
As illustrated in Figure 2, the xGen-MM (BLIP-3) framework adopts an architecture consisting of a ViT [44, 45], a vision token sampler (perceiver resampler [22]) to downsample the image embeddings, and a pre-trained Large Language Model (phi3-mini [37]). The input to the model can be free-form multimodal interleaved texts and vision tokens from the diverse multimodal data sources we ensemble.
Any-Resolution Vision Token Sampling.
As proved effective in recent LMMs [46, 47, 48], we adopt a dynamic high-resolution (i.e., "any-resolution") image encoding strategy at the fine-tuning and post-training stages. We enable higher-resolution image understanding with image patch-wise encoding. The patch-wise encoding preserves the resolution of the original image as much as possible by splitting a single image into multiple patches and encoding them separately. Following the previous convention, we concatenate the encoded image patches with a downsized original image that provides global information.
@@ 4 Training
Pre-training.
The pre-training objective is to predict the next text token across the dataset mixture we pre-train on. Overall, the resulting base model xGen-MM-Phi3-mini-base-r is pre-trained for about 100 billion multimodal tokens from the ensembled dataset, and our pre-training resolution is 384x384 pixels, which aligns with SigLIP [45].
Supervised Fine-tuning (SFT).
We further fine-tune our pre-trained models on instruction-following examples to make them better understand and follow user queries. At the fine-tuning stage, we use a collection of publically available instruction-following datasets [11, 29, 49]. We adopt the any-resolution vision token sampling strategy to allow a better understanding of images of higher resolutions such as text-rich document data. We introduce the technical details for the fine-tuning stage in the following sections.
Interleaved Multi-Image Supervised Fine-tuning.
We conduct a second-stage fine-tuning on the instruction fine-tuned model on a mixture of multi-image and single-image instructions-following samples. The goal for this second-stage fine-tuning is to enhance the model’s ability to comprehend interleaved image-text input, which is helpful for multimodal in-context learning, multi-image question answering, and many more practical use cases. For the multi-image fine-tuning, we also adopt the any-resolution vision token sampling strategy same as in the previous SFT stage.
Post-training.
Finally, we perform two stages of post-training to improve the model’s helpfulness while mitigating harmful qualities such as hallucination and toxicity. We first perform direct preference optimization (DPO [50]) to improve the model’s helpfulness and visual faithfulness. We then perform safety fine-tuning to improve the model’s harmlessness. We quantitatively demonstrate Pareto gains in model harmlessness and helpfulness after our post-training.
@@ 5.1 Pre-training Data Recipe
As indicated in Figure 3, in xGen-MM (BLIP-3), we pre-train on an ensemble of diverse multimodal datasets with the indicated sampling ratios.
@@
Figure 3: Overview of xGen-MM (BLIP-3) Pre-training Datasets.
Interleaved Dataset Mixture.
@@
We combine MINT-1T (including its HTML, PDF, and ArXiv subsets) with OBELICS (HTML only) to create a more diverse and comprehensive dataset mixture that covers a broader range of domains.
  1. 1.
@@
    MINT-1T [12] is a trillion token scale multimodal interleaved dataset, containing data sources from HTML, PDF, and ArXiv. As evidenced by MM1 [9] and Idefics2 [11], such multimodal interleaved datasets are essential for scaling up large multimodal model training and enabling fundamental capabilities like multimodal in-context learning. Notably, different from the OBELICS [11], MINT-1T has three subsets from different sources: the HTML subset, the PDF subset, and the ArXiv subset. In xGen-MM (BLIP-3), these three subsets are mixed in a 7:5:1 ratio.
  2. 2.
@@
    OBELICS [11] is another large-scale multimodal interleaved dataset constructed from HTML documents solely. It differs slightly in domain coverage from MINT-1T due to the specific preprocessing steps adopted.
Caption Dataset Mixture.
@@
We integrate a diverse range of caption datasets, with the specifics outlined in the following details.
  1. 1.
    BLIP3-KALE is a large-scale curated high-quality caption dataset. Details will be discussed in another paper, and the dataset will be made public very soon.
  2. 2.
    BLIP3-OCR-200M is a curated large-scale OCR dataset to address the limitations of current large multimodal models in handling text-rich images like documents and charts, as traditional image-text datasets often lack adequate OCR annotations. To enhance text comprehension abilities, we use the OCR engine PaddleOCR [51] to annotate images with OCR-specific annotations. Overall, we curate a dataset of 200 million high-resolution images from Datacomp-1B[17]. For each image, we create captions with OCR data by identifying and extracting textual elements using the off-the-shelf OCR engine [51]. Text segments in a caption like "… text …" are modified to include OCR information as "… text ( ocr_info ) …", where ocr_info contains bounding box coordinates for the extracted text, specifying its exact position within the image in the format "<bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>". We have multiple granularities of OCR information, including with and without bounding box data. In our work, we only utilize textual information without bounding box data, which has proven to be effective. Note that, in xGen-MM (BLIP-3), we preprocess the captions to remove filler texts such as “the text”, which we find improves OCR-related benchmark performance. We hypothesize that this is because such filler text is relatively easy to predict, potentially diluting the loss associated with OCR-relevant tokens. Nonetheless, incorporating bounding box information could further enhance performance, and we encourage researchers in the community to explore this potential.
  3. 3.
    BLIP3-GROUNDING-50M is a curated large-scale grounding dataset to enhance the ability to ground semantic concepts in visual features, which is crucial for tasks like object detection, semantic segmentation, and understanding referring expressions [52] (e.g., "the object to the left of the dog"). We curate a dataset of 50 million images from Datacomp-1B [17]. For each image, we identify objects and their location information using one of the state-of-the-art open-world image tagging [53] and object detection models [54]. Objects mentioned in a caption like "… object …" are modified to include grounding information as " … object ( grounding_info ) …", where grounding_info contains bounding box information in one of three formats, each capturing a different granularity of localization: (1) <bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>, (2) "starts at $(x_{1},y_{1})$ and extends up to $(x_{2},y_{2})$", or (3) "top-left corner of the image".
  4. 4.
    Other Public Datasets Mixture: We also include other publicly available datasets such as uncurated Datacomp-1B [17] image-text pairs, CC12M [14], CC3M [14], VG [15], and SBU [16].
Figure 4: Samples from BLIP3-OCR-200M. Six levels of OCR information granularity are extracted, with and without bounding box data. Note that in xGen-MM (BLIP-3), OCR-related captions are preprocessed to remove filler phrases like ’the text,’ resulting in improved OCR benchmark performance.
@@
Figure 5: Samples from BLIP3-GROUNDING-50M. We introduce a large-scale dataset of images and corresponding captions containing localization information about objects. Furthermore, we release the associated object bounding box data to facilitate the creation of captions with custom templates.

### v3

# BLIP-3: A Family of Open Large Multimodal Models

@@ Abstract
This paper introduces BLIP-3, an open framework for developing Large Multimodal Models (LMMs). The framework comprises meticulously curated datasets, a training recipe, model architectures, and a resulting suite of LMMs. We release 4B and 14B models, including both the pre-trained base model and the instruction fine-tuned ones. Our models undergo rigorous evaluation across a range of tasks, including both single and multi-image benchmarks. Our models demonstrate competitive performance among open-source LMMs with similar model sizes. Our resulting LMMs demonstrate competitive performance among open-source LMMs with similar model sizes, with the ability to comprehend interleaved image-text inputs. Our training code, models, and all datasets used in this work, including the three large-scale datasets we create and the preprocessed ones, will be open-sourced to better support the research community. Refer to caption Figure 1: We introduce BLIP-3, a framework (b) for developing Large Multimodal Models (LMMs). Our framework improves upon BLIP-2 (a) [42] by (1) increasing the richness, scale, and diversity of training data, (2) replacing the Q-Former layers with a more scalable vision token sampler, and (3) simplifying the training process via the unification of the training objectives to a single loss at every training stage. (c) illustrates the stage-by-stage training process, from initializing with a pre-trained LLM to the final LMMs.
@@ 1 Introduction
Large Multimodal Models (LMMs) have attracted significant attention with their potential applications and emergent capabilities. Recent advancements in both proprietary models [63, 25, 64, 82] and open-source LMMs [41, 42, 15, 49, 60, 47, 38, 29, 46, 2, 26, 55, 71, 66] highlight the rapid progress and growing interest in this field. However, despite these advancements, a significant gap remains between open-source and proprietary models, particularly in terms of accessibility to good base models, training recipes, and curated datasets. While there are many strong open-weight models [1, 2, 26], not many of them disclose their training details or release their full training data recipe (including training code). This lack of transparency limits the broader research community’s ability to replicate, analyze, and improve LMMs. Additionally, many of these models do not contribute new datasets, further restricting progress. These constraints hinder transparency and innovation, preventing open-source communities from fully leveraging and advancing LMM research.
@@
Recent works have demonstrated that large-scale and high-quality data are essential for training robust LMMs [49, 60, 47, 38, 6, 7]. BLIP-2 [42] was one of the pioneering efforts in exploring LMMs, which leveraged synthetic data to achieve impressive results at the time (Figure 1 (a)). However, the data used in BLIP-2 lacks the scale, quality, and diversity required to reach competitive performance compared to more modern LMMs nowadays. In addition, BLIP-2 employs an intricate Q-Former [42] architecture to bridge the vision and language modalities, coupled with a suite of complex training objectives (ITM, ITC, and ITG losses), both of which pose obstacles for larger-scale training. Moreover, BLIP-2 supports only single-image input, whereas interleaved multimodal data formats are the most natural form of multimodal data [37].
@@
In response to these challenges, we introduce BLIP-3 (Figure 1 (b)), a new framework designed to scale up LMM training by utilizing an ensemble of multimodal interleaved datasets, curated caption datasets, and other publicly available datasets [11, 36, 65, 22, 9]. In BLIP-3, as illustrated in Figure 2, we streamline the model architecture by replacing the Q-Former [42] with a more scalable vision token sampler [4] and simplifying the training objectives to focus solely on the auto-regressive loss of text tokens in a multimodal context. Our primary focus is on dataset curation and scaling up the training data. Recently, several high-quality, large-scale datasets have been introduced, such as MINT-1T[6], a trillion-token-scale interleaved dataset, and BLIP3-KALE[7], a knowledge-augmented dataset with high-quality dense captions. In this paper, we introduce three additional specialized datasets: BLIP3-OCR-200M, a large-scale dataset with dense OCR annotations for general base-resolution OCR understanding; BLIP3-GROUNDING-50M, a large-scale visual grounding dataset; and BLIP3-OCR-HD-30M, a high-resolution dataset with high-quality OCR annotations, designed for high-resolution pre-training.
@@
In addition to these datasets, we are committed to open-sourcing the series of models developed in this work, including both the pre-trained base models and instruction-tuned models. Along with the model release, we also provide our training code. By making these resources publicly available, we aim to make LMM research and development more accessible to the community, and we encourage researchers and practitioners to use our models and datasets to understand and further explore the potential and emergent capabilities of LMMs.
@@
Figure 2: Overview of the BLIP-3 architecture. Free-form interleaved images and texts from the ensembled interleaved and caption datasets are input into the framework, with each modality undergoing a separate tokenization process to be fed into the pre-trained LLM in natural order. A standard auto-regressive loss is then applied to the text tokens. The Vision Transformer is kept frozen during training, while all other parameters, including the vision token sampler and the pre-trained LLM, are trained.
@@ 2 Related Work
Recent advancements in Large Multimodal Models (LMMs) have explored different architectural designs. We roughly divide them into two main categories: the cross-attention style [4, 5] and the self-attention style. The cross-attention approach, exemplified by models like Flamingo [4, 5] and Llama 3.1 [55], integrates vision and language modalities by injecting visual information into layers inside LLMs using a cross-modal cross-attention module. The self-attention approach [42, 15, 49, 8, 14, 17, 30, 67, 48, 77, 78, 24, 10, 23, 70], on the other hand, offers a more streamlined solution. This approach connects pre-trained language models to visual inputs using lightweight connectors. Both visual and textual inputs are tokenized and concatenated in the same input sequence to LLM, and the LLM learns to align both modalities through self-attention. The self-attention design has become prevalent recently, which has been shown to be effective on a wide range of LMMs [50, 1, 2, 76, 8] . For BLIP-3, we also adopt the self-attention design due to its simplicity.
Training methodologies for LMMs typically follow one of the two strategies. The first one uses a light pre-training procedure and heavily relies on visual instruction tuning, as seen in the LLaVA series [49, 48]. Extensive research has been conducted on creating effective instruction-tuning data for a variety of tasks [24, 12, 43, 86, 73, 40]. The second strategy involves extensive pre-training on large-scale, diverse datasets, followed by visual instruction fine-tuning. This approach, used by models such as MM1 series [60, 82] and Idefics2 [38], infuses a broad knowledge into the model during the pre-training stage, and then fine-tune it for better instruction following ability and better alignment. While MM1 series [60, 82] provides extensive ablations and studies on the recipes aimed at improving LMMs, it releases limited resources for practitioners to reproduce the model (their models and datasets are not open-sourced).
In this work, we introduce BLIP-3, an open LMM family encompassing a series of models, data recipes, training code, and three new large-scale foundational multimodal datasets. Unlike previous works, BLIP-3 is designed to enable and advance future research in this area by providing comprehensive resources for the community.
@@ 3 Model Architecture
As illustrated in Figure 2, the BLIP-3 framework adopts an architecture consisting of a ViT [19, 81], a vision token sampler (perceiver resampler [4]) to downsample the image embeddings, and a pre-trained LLM (Phi-3 series [1]). The input to the model can be free-form multimodal interleaved texts and vision tokens from the diverse multimodal data sources we ensemble.
Any-Resolution Vision Token Sampling. As proved effective in recent LMMs [50, 83, 16], we adopt a dynamic high-resolution (i.e., “any-resolution”) image encoding strategy at both the Stage-2 pre-training and the fine-tuning stages. We enable higher-resolution image understanding with image patch-wise encoding. The patch-wise encoding preserves the resolution of the original image as much as possible by splitting a single image into multiple patches and encoding them separately. Following the previous convention, we concatenate the encoded image patches with a downsized original image that provides global information.
@@ 4 Training
Stage-1 Base Resolution Pre-training. The first stage of pre-training focuses on base resolution image-text alignment using a diverse dataset mixture comprising open-source datasets and two newly created datasets from BLIP-3 – BLIP3-OCR-200M and BLIP3-GROUNDING-50M, all of which will be made publicly available. In this stage, the base model is trained on approximately 100 billion multimodal tokens from the ensembled dataset. To ensure computational efficiency, the pre-training resolution is set to 384×384 pixels, aligning with the default resolution of SigLIP [81].
Stage-2 High Resolution Pre-training. Building upon the pre-trained checkpoint from Stage 1, the second stage incorporates additional high-quality, high-resolution datasets, including one newly created dataset from BLIP-3 – BLIP3-OCR-HD-30M, to further enhance the model’s capability. To maintain consistency with the supervised fine-tuning (SFT) stage, we adopt the same any-resolution vision token sampling strategy introduced in section 3. We set the any-resolution grid to support up to 12 image patches with a maximum side length of 1536 pixels on the longer side. Further details and ablation studies are provided in section 7.2.
Single-Image Supervised Fine-tuning. Then, we fine-tune the pre-trained model on a collection of publicly available instruction-following datasets [38, 48, 71]. During fine-tuning, we employ the same any-resolution vision token sampling strategy as in Stage 2, enabling the model to process high-resolution images effectively, including text-rich document-style data. The following sections provide additional technical details on the fine-tuning procedure.
Interleaved Multi-Image Supervised Fine-tuning. We conduct a second-stage fine-tuning on the instruction fine-tuned model on a mixture of multi-image and single-image instruction-following samples. The goal of this second-stage fine-tuning is to enhance the model’s ability to comprehend interleaved image-text input, which is helpful for multimodal in-context learning, multi-image question answering, and many more practical use cases. For the multi-image fine-tuning, we also adopt the any-resolution vision token sampling strategy, the same as in the previous SFT stage.
@@ 5.1 Pre-training Data Recipe
Our Stage-1 pre-training recipe is detailed in Figure 3. In BLIP-3, we pre-train on a diverse set of multimodal datasets using the specified sampling ratios.
For Stage-2 pre-training, we use the Stage-1 data recipe as a base while incorporating specialized high-resolution, text-rich datasets. Specifically, we replace BLIP3-OCR-200M with BLIP3-OCR-HD-30M and introduce the IDL [9] dataset, each making up 10% of the total training data. To preserve the overall structure of the dataset mixture, we adjust the sampling ratios of the remaining caption datasets by reducing BLIP3-KALE from 25% to 10%, Datacomp-1B from 10% to 2.5%, and BLIP3-GROUNDING-50M from 5% to 2.5%, while keeping the rest unchanged.
@@
Figure 3: Overview of BLIP-3 Stage-1 pre-training datasets base recipe.
@@
Interleaved Dataset Mixture. We combine MINT-1T (including its HTML, PDF, and ArXiv subsets) with OBELICS (HTML only) to create a more diverse and comprehensive dataset mixture that covers a broader range of domains.
1. MINT-1T [6].
@@
A trillion-token-scale multimodal interleaved dataset, containing data sources from HTML, PDF, and ArXiv. As evidenced by MM1 [60] and Idefics2 [38], such multimodal interleaved datasets are essential for scaling up large multimodal model training and enabling fundamental capabilities like multimodal in-context learning. In BLIP-3 we mix its HTML, PDF, and ArXiv subsets in a 7:5:1 ratio.
2. OBELICS [38].
@@
A multimodal interleaved dataset constructed solely from HTML documents. It differs slightly in domain coverage from MINT-1T due to the specific preprocessing steps adopted.
@@
Caption Dataset Mixture. We integrate a diverse range of caption datasets, with the following details.
@@
Figure 4: Samples from BLIP3-OCR-200M. We extract six levels of OCR granularity, with and without bounding boxes. OCR-related captions in BLIP-3 are preprocessed to remove filler phrases like ’the text’ within the dataloader, which we find improving OCR benchmarks’ performance. Samples from BLIP3-OCR-HD-30M. Structured similarly to BLIP3-OCR-200M, this dataset contains high-resolution images annotated with high-resolution OCR tools for enhanced high resolution OCR understanding. Only Level 1 is shown for simplicity. Samples from BLIP3-GROUNDING-50M. A large-scale dataset of images and corresponding captions containing localization information about objects. Furthermore, we release the associated object bounding box data to facilitate the creation of captions with custom templates.
1. BLIP3-KALE [7].
An open-source large-scale curated high-quality caption dataset, which we use as a base for general detailed caption datasets.
2. BLIP3-OCR-200M.
A curated large-scale OCR dataset with 200 million images from Datacomp- 1B[22]. For each image, we use the off-the-shelf OCR engine [28] for OCR annotation. Text segments in a caption like "… text …" are modified to include OCR information as "… text ( ocr_info ) …", where ocr_info contains normalized bounding box coordinates for the extracted text, specifying its exact position within the image in the format "<bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>". We have multiple granularities of OCR information, including with and without bounding box data. In BLIP-3 pre-training, we only utilize textual information without bounding box data.
3. BLIP3-GROUNDING-50M.
A curated large-scale grounding dataset to enhance the ability to ground semantic concepts in visual features, which is crucial for tasks like understanding referring expressions [79] (e.g., "the object to the left of the dog"). We curate a dataset of 50 million images from Datacomp-1B [22]. For each image, we identify objects and their location information using one of the state-of-the-art open-world image tagging [85] and object detection models [51]. Objects mentioned in a caption like "… object …" are modified to include grounding information as " … object ( grounding_info ) …", where grounding_info contains bounding box information in one of three formats, each capturing a different granularity of localization: (1) <bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>, (2) "starts at $(x_{1},y_{1})$ and extends up to $(x_{2},y_{2})$", or (3) "top-left corner of the image".
4. BLIP3-OCR-HD-30M.
A high-resolution OCR dataset curated for Stage-2 pre-training. Different from BLIP3-OCR-200M, which focuses on more general OCR data with fewer resolution constraints, this dataset is filtered to only contain images of which both the height and width are larger than 512 pixels. On the selected high-resolution images, we run PaddleOCR with a high-resolution setup to obtain more accurate OCR annotations. Our empirical study shows that running the OCR annotation process with a high-res setting can help get a more complete and accurate set of OCR annotations that matches its resolution, which is more useful for the Stage-2 pre-training. This process yields a total of 30 million high-resolution samples with detailed OCR annotations that match the image quality.
5.Other Datasets Mixture.
We also include other publicly available datasets such as uncurated Datacomp-1B [22] image-text pairs, CC12M [11], CC3M [11], VG [36], SBU [65], and IDL [9].